1. Doctor orders digoxin 0.92 mg daily for a child that weighs 16 lbs. The safe dosage for this medication is 8-12 mcg/kg/day. Is this a safe dosage for this child/ explain

2. Doctor orders Amoxicillin 500mg PO BID times 10 days for infection. Available is Amoxicillin 250 mg/10ml.

How many ml should the nurse give

3. A child weighs 100 lbs.The child has a bacteria infection and the doctor orders penicillin I gram stat. Available penicillin is 500 mg/ml. Ho many ml should the nurse give?

4. Order: Acetaminophen elixir 15 mg/kg po. Child weight is 15 pounds. Available Acetaminophen elixir is 100 mg/5ml. How many ml will you administer?
Answer #1

ANSWER.

1) Safe dose for the child.

         Ordered dose = 0.92 mg.

         Weight of the child = 16 lbs .

    Convert lbs to kg for calculation.

               1 lbs = 0.45 kg.

      So

         Weight of the child = 16 * 0.45

                                        = 7.2 kg.

       Here

          Safe dosage for this medication = 8-12 mcg/kg/day.

                  Convert mcg value to mg for calculation

                                 1 mcg = 0.001 mg.

                     So

                       Safe dose for this medication = 0.008 mg to 0.012 mg.

      Safe dose for this child with 7.2 kg body weight = 0.008 * 7.2 - 0.012 * 7.2

                                                                  = 0.06mg to 0.09mg / day.

       * Ordered dose (0.92) is more than safe dose so it is not safe for this child.

2) Calculation of amount(ml) of drug required.

          Formula = Desired dose/Dose in hand *Quantity.

     Here

          Desired dose(D) = 500 mg.

          Dose in hand (H) = 250 mg.

                 Quantity (Q) = 10 ml.

As per formula

      Amount(ml) of drug required = 500/250 * 10

                                                     = 2 * 10

                                                     = 20 ml.

3) Calculation of amount(ml) of drug required.

           Formula = Desired dose /Dose in hand * Quantity.

       Here

            Desired dose (D) = 1 gm.

         Convert gm value to mg for calculation

                       1gm = 1000 mg.

           So

               Desired dose(D) = 1000 mg.

               Dose in hand (H) = 500 mg.

                  Quantity(Q) = 1 ml.

    As per formula

         Amount(ml) of drug required = D/H * Q

                                                        = 1000/500 * 1

                                                        = 2 * 1

                                                        = 2 ml.

4) Calculation of amount of drug required.

          Formula = Desired dose /Dose in hand * Quantity.

     Here

         Ordered dose = 15 mg/kg.

         Child weight = 15 lbs.

       Convert lbs value to kg for calculation

                  1 lb = 0.45 kg.

          Child weight = 15 * 0.45

                               = 6.8 kg.

          Desired dose (D) = 15 * 6.8

                                       = 102 mg.

          Dose in hand (H) = 100 mg.

                   Quantity(Q) = 5 ml.

   As per formula

         Amount(ml) of drug required = D/H * Q

                                                       = 102/100 * 5

                                                       = 1.02 * 5

                                                       = 5.1 ml.